OpenStudy (anonymous):

sin(2460°)? I am getting answer as -root3/2??

OpenStudy (neer2890):

correct...!

OpenStudy (anonymous):

yuupp..its right ans..

OpenStudy (kropot72):

\[\sin 2460=\sin -60=-\frac{\sqrt{3}}{2}\]